# Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge

### v1

@@ 4.3 Probing Metric
Both mAP and mWS metrics range from [0,1].
5 Experiments
5.1 Access evaluation
We explore the masked language models, including the encoder-only language model BERT Devlin et al. (2019) and its multilingual version, mBERT, along with the encoder-decoder language model T5 Raffel et al. (2020) and its multilingual counterpart, mT5 Xue et al. (2020). Table 2 showcase the probing results based on English prompt on the filtered dataset FmLAMA-en. [^5] The probing results with prompts in the other five languages on the corresponding filtered sub-datasets can be found in the Appendix C.
Metric comparison
Regarding the adopted two metrics, we observe the following: 1) Across diverse cultural groups, evaluating probing results with the absolute-match metric, mAP, reveals a significantly greater disparity compared to assessments using the fuzzy-match metric, mWS. This highlights the challenge of absolute-match evaluation and the importance of introducing the fuzzy-match metric. 2) We evaluate the correlation between the two metrics, mAP, and mWS, utilizing all probing results from Table 2, resulting in a Pearson correlation coefficient of 0.66. This signifies a moderate to strong positive correlation between the two metrics.
Group comparison
Regarding the various cultural groups in Table 2, we observe the following: (1) Regardless of absolute or fuzzy evaluation, the U.S. group almost consistently achieves the highest probing results on FmLAMA-en with English prompts. This suggests that evaluated LLMs possess a greater familiarity with food-related knowledge specific to the U.S. context. (2) The quantity of knowledge within cultural groups in the knowledge base may not fully reflect the potential knowledge within LLMs. For example, despite the Italy group having more dishes than the U.K. and India groups, LLMs achieve lower probing result scores for the Italy group. We speculate that this difference may be due to the official languages of the U.K. and India including English, aligning with the prompt language in the experiments. However, it is notable that LLMs do not exhibit lower probing result scores in China, despite its official language not including English. (3) Monolingual English LLMs excel in accessing cultural knowledge within English-speaking countries, while multilingual LLMs may not demonstrate superior performance in non-English-speaking countries. This discrepancy may arise from the fact that the pretraining data for monolingual models is solely in English, whereas for multilingual models, it encompasses languages beyond English.
@@ 5.1 Access evaluation
Figure 2: Comparative probing results on FmLAMA-en: incorporating cultural information about the origin of dishes into English prompts can enhance the probing of cultural knowledge.
5.2 Prompt analysis
@@
In this part, we examine the impact of different prompt settings on cross-cultural knowledge exploration. This involves integrating references to cultural backgrounds, incorporating code-switching settings, and considering language choices within the prompts.
Cultural background analysis
@@
Using the English prompts from Figure 7 as a basis, we incorporate information about the country of origin into the probing prompts for each specific dish (as described in §4.1). Figure 2 presents a comparative analysis of probing results on FmLAMA-en across different LLMs, taking into account the inclusion of cultural information mentions in the prompts. We find that, whether through absolute or fuzzy evaluation, English prompts with cultural information achieve higher probing scores in capturing LLMs’ knowledge within the food domain. This suggests the importance of emphasizing the cultural background when utilizing LLMs, especially in the exploration of culture-related topics. Moreover, in the comparison of the absolute-match metric mAP, cultural information contributes more in multilingual LLMs compared to monolingual LLMs. This seems to align with real-world social dynamics, indicating that in multicultural settings, introducing cultural backgrounds enhances communication.
Figure 3: Probing results with monolingual prompts and code-switching prompts, $\circ$ and $\Box$ signify the absence or presence of cultural background information in the prompt, respectively.
Code-switching analysis
Code-switching probing focused on dishes common in both English and Chinese in a filtered subset of FmLAMA. Only cultural groups with dish counts exceeding 20 are retained. Our objective is to ensure that the language of the predicted ingredients matches the main language used in the prompt. Probing results on each multilingual LLMs display for Figure 3: (1) In code-switching settings, the probing abilities of LLMs for EN_zh relative to EN_en and ZH_en relative to ZH_zh both notably decrease, with the latter showing a more pronounced decline in the mB-c language model. (2) Whether in code-switching mode or not, the introduction of relevant cultural backgrounds in the prompts generally facilitates cultural knowledge probing. The detailed probing results of each cultural group can be seen in Appendix D.
Figure 4: Average probing results of all multilingual LLMs on the filtered sub-dataset with prompts in different languages. The ability of LLMs to access cultural knowledge may not consistently outperform in the cultural group where that language is spoken for each language prompt.
Language analysis
We adopt prompts in six different languages and conduct knowledge probing on multilingual LLMs. To ensure a fair comparison of the probing results across different language prompts, we filter a sub-dataset consisting of only 175 food instances that have labels in all of the languages that are involved. This underscores the lack of aligned knowledge across multiple languages. Considering the relatively limited size of the filtered dataset, we opt for a broader categorization based on continents rather than individual countries to differentiate cultural groups. Appendix E illustrates the distribution of cultural groups within this filtered sub-dataset. Figure 4 displays probing results on the filtered sub-dataset with prompts in different languages. The probing performance varies with prompts in different languages. The ability of LLMs to access cultural knowledge may not consistently outperform in the cultural group where that language is spoken for each language prompt. For instance, Chinese prompts yield better results in English-speaking countries. Our experimental results suggest that determining the most suitable language for knowledge probing in a specific cultural context proves challenging.
Prompt language EN ZH
Subject language en zh en zh
Llama2-7b-chat-hf 27.36 17.91 7.28 5.91
Llama2-13b-chat-hf 35.04 22.64 10.43 7.87
vicuna-7b 36.22 19.69 12.01 14.37
vicuna-13b 28.15 20.87 17.72 16.54
gpt-3.5-turbo 47.64 34.06 20.28 25.00
Table 3: Accuracy (%) of decoder-only LLMs’ results.
5.3 Decoder-only LLMs probing
LLMs, particularly decoder-only models, offer robust capabilities in both natural language understanding and generation. We specifically conduct cultural knowledge probing on instruction-tuned models, including Llama2-chat Touvron et al. (2023), Vicuna Chiang et al. (2023), and gpt-3.5-turbo OpenAI (2023). Considering the characteristics of these LLMs, we employ a naive approach to probe and evaluate them. [^6] The experimental setup details, evaluation method, cultural knowledge probing results for each cultural group, and additional analysis are provided in Appendix F. Table 3 shows the overall probing results using accuracy for both the monolingual and code-switching prompt settings. We observe that gpt-3.5-turbo demonstrates the highest extent of cultural knowledge. Additionally, all decoder-only LLMs demonstrate better performance with monolingual English prompts compared to monolingual Chinese prompts, and the performance with code-switching prompts is inferior to that with their corresponding monolingual prompts.
Most common ingredients Gold label ingredients
en zh ko en zh ko
yogurt water tangerine egg grain flour egg
avocado sugar cumin seeds flour egg wheat flour
vegetable glucose hot stone sugar sugar sugar
olive oil oil lime juice potato rice potato
rice salt pork floss almond chicken salt
bread egg rice meat tomato chicken meat
baking powder quince drinking water onion onion butter
ice cream milk apple tomato butter meat
extra virgin olive oil sugar cheese chicken meat pork milk
vegetable oil cilantro tofu butter potato bread
Table 4: Ingredients analysis. The first three columns display the 10 most common ingredients for mBERT-base-uncased in English (en), Chinese (zh), and Korean (ko). The remaining three columns show the 10 most common gold standard ingredients, sourced from Wikipedia.
5.4 Case study – Ingredients analysis
We conduct a more fine-grained analysis of the model’s predictions to better understand its behavior, particularly to discern whether it is relying on “educated” guesses or leveraging its cultural knowledge when providing answers. We focus on two multilingual models, mT5 and mbert-base-uncased, and three languages: Korean, Chinese, and English [^7] This subset of languages is selected for clarity. The observed trends are consistent across all languages in our dataset.. For every language and dish in the multilingual dataset (see §3), we extract the top 5 ingredients with the highest probability as predicted by the model and calculate the set of distinct ingredients. Table 4 shows mbert’s 10 most common ingredients in each language. In Korean, Chinese, and English, these ingredients cover $0.59\%$, $0.91\%$, and $0.97\%$ of the total predictions, respectively. [^8] To compute the coverage we divide the number of times the most common 10 predictions are predicted by the model, by the number of its top 5 predictions. For each dish we take only 5 predictions because more than $0.95\%$ of the dishes contain less than 5 ingredients. This indicates that the model tends to depend on repetitive predictions, consistently selecting the same subset of ingredients regardless of the context.  [^9] This trend persist for all models, and languages. For example, in Korean, mbert often predicts ingredients specific to Korean dishes, such as hot stone, which refers to various Korean dishes served in a hot stone pot, and tofu. However, in Chinese/English, the model tends to predict more widely used ingredients, like flour and sugar, which are common across various cultural cuisines. This observation might imply that Chinese/English models perform better because the ingredients they predict have broader applicability and are more likely to be found in a variety of dishes.
6 Human evaluation
To check the validity of the two proposed metrics, we conduct a human evaluation of the probing results, including 372 food instances. Specifically, we extract the probing results for all dishes originated in the U.S and China from the best-performing prompt setting in the code-switching EN_en experiments. In each dish instance, only the top-$l$ predicted objects are evaluated, where $l$ is the number of golden labels, similar to the setting of the mWS metric. In addition to assessing the absolute matching between predicted objects and golden objects as measured by mAP, our human evaluation also considers: (1) whether a predicted object can replace a certain golden object in cooking, and (2) whether the predicted object is an ingredient of the dish, even if not listed in the golden label (e.g, lemon is an ingredient of the dish lemon chicken, even though it erroneously has only chicken meat in Wikidata).
The calculation method for human evaluation scores is the same as mAP. Four authors participated in human evaluation for this work, who come from three different countries, three of whom have experience living across various cultures. The Krippendorff’s alpha value about inter-annotator agreement is 0.581, which suggests a moderate level of agreement among raters. We average their final scores to obtain the ultimate assessment score. Table 5 compares the scores from human evaluation with those from two automated metric evaluations. The Pearson correlation coefficient between mAP and Human is 0.88, whereas that between mWS and Human is 0.94 as per the aforementioned results, which indicates the significance of mWS, as it aligns more closely with human evaluation.
Origin mBERT-uncased mT5
mAP mWS Human mAP mWS Human
U.S. 10.68 37.96 33.61 14.70 38.24 37.56
China 8.40 32.79 25.16 9.40 37.41 32.19
Table 5: Comparison of scores (%) from human evaluation and two automated metric evaluations. The Pearson correlation coefficient between mAP and human is 0.88, whereas that between mWS and human is 0.94 as per the aforementioned results.

### v3

@@ 4.3 Probing Metric
Encoder-only LLMs Encoder-Decoder LLMs Decoder-only LLMs
Origin Count Dumb Base Bb Bl mB T5 mT5 Qwen2 Llama2 Llama3 Avg.
Italy 215 (13.9%) 18.14 8.12±1.14 7.52±0.43 9.31±1.33 4.54±0.56 6.73±1.34 26.82±4.36 19.37±3.68 29.62±3.30 14.00
U.S. 285 (18.4%) 11.10 19.98±2.00 18.83±1.95 19.91±5.88 10.24±2.62 16.50±4.86 32.21±2.05 21.64±5.75 36.58±3.58 21.99
Turkey 98 (6.3%) 12.90 11.80±1.03 9.84±3.81 14.98±3.89 6.10±1.57 10.25±2.91 25.22±1.46 22.96±4.04 30.01±2.30 16.39
Japan 186 (12.0%) 9.35 11.10±1.19 9.14±1.81 11.52±1.89 5.45±1.90 6.38±1.00 23.15±2.01 20.24±3.10 25.47±1.73 14.06
France 175 (11.3%) 16.50 9.33±1.24 7.51±0.90 8.80±2.34 3.86±1.62 4.57±1.11 29.56±2.38 21.95±4.83 30.01±2.19 14.45
U.K. 83 (5.4%) 18.67 12.47±1.77 13.72±2.20 14.20±4.01 6.16±2.89 9.88±3.06 29.06±2.47 21.82±3.43 32.15±4.06 17.43
Mexico 57 (3.7%) 9.40 8.25±1.04 10.35±2.88 8.87±2.18 3.41±0.49 5.96±2.36 25.14±1.58 22.03±5.65 28.97±3.54 14.12
India 132 (8.5%) 11.57 18.42±1.35 16.94±3.10 17.51±1.75 9.18±2.98 7.80±1.15 35.82±1.79 33.88±5.00 37.62±1.89 22.15
Germany 57 (3.7%) 14.02 8.54±1.77 7.84±1.01 9.00±2.39 4.38±1.48 6.68±1.56 29.05±2.63 20.42±5.56 28.86±3.78 14.35
China 97 (6.3%) 8.60 15.06±2.45 14.94±1.37 15.24±5.31 7.80±1.97 12.39±2.62 28.96±3.15 21.02±4.78 30.39±2.70 18.23
Iran 21 (1.4%) 12.60 9.16±1.75 8.19±1.51 13.56±3.16 9.00±2.04 6.21±0.15 40.26±8.19 35.36±9.81 45.72±3.65 20.93
Greece 21 (1.4%) 15.07 4.09±1.10 3.48±1.45 4.85±1.37 3.81±1.09 1.32±0.15 31.68±8.85 15.88±2.54 31.60±3.17 12.09
Spain 95 (6.1%) 16.05 9.18±0.93 7.18±1.04 7.42±1.51 3.98±1.37 3.33±0.46 25.21±5.07 17.00±4.23 25.36±1.68 12.33
Russia 27 (1.7%) 10.64 4.72±0.98 7.10±2.81 6.44±1.97 1.41±0.60 2.21±0.33 14.83±4.80 10.47±0.98 11.77±0.59 7.37
ALL 1549 (100.0%) 13.29 12.56±1.06 11.58±1.29 12.96±2.84 6.32±1.42 8.68±2.03 28.56±2.27 21.89±4.06 30.96±2.09 16.69
Coefficient of Variation (CV) 24.15 40.87 40.76 36.89 43.29 54.29 20.39 28.27 23.84 -
(a) Performance results evaluated using mAP (%). CV indicates the extent of cultural bias in each model’s performance.
Origin Encoder-only LLMs Encoder-Decoder LLMs Decoder-only LLMs Avg.
Bb Bl mB T5 mT5 Qwen2 Llama2 Llama3
Italy 0.3813±0.02 0.3624±0.01 0.3573±0.02 0.3129±0.02 0.3195±0.01 0.4930±0.02 0.4510±0.02 0.5057±0.01 0.3979
U.S. 0.4461±0.02 0.4358±0.02 0.4219±0.06 0.3516±0.02 0.4076±0.03 0.5147±0.02 0.4656±0.04 0.5411±0.02 0.4481
Turkey 0.4130±0.02 0.3886±0.03 0.3954±0.04 0.3387±0.03 0.3419±0.02 0.4689±0.01 0.4612±0.03 0.5022±0.01 0.4137
Japan 0.3778±0.02 0.3501±0.03 0.3446±0.03 0.3059±0.03 0.2932±0.01 0.4416±0.01 0.4304±0.02 0.4436±0.01 0.3734
France 0.3750±0.02 0.3660±0.02 0.3497±0.04 0.3279±0.02 0.3367±0.01 0.5201±0.02 0.4697±0.03 0.5099±0.01 0.4069
U.K. 0.3708±0.02 0.3775±0.02 0.3583±0.04 0.3031±0.02 0.3320±0.03 0.4532±0.03 0.4291±0.01 0.4620±0.04 0.3857
Mexico 0.3509±0.01 0.3654±0.03 0.3255±0.03 0.3287±0.01 0.3262±0.02 0.4547±0.02 0.4762±0.04 0.4863±0.03 0.3892
India 0.4402±0.01 0.4193±0.04 0.4135±0.02 0.3380±0.04 0.3141±0.01 0.5106±0.00 0.5140±0.03 0.5190±0.02 0.4336
Germany 0.3699±0.03 0.3492±0.03 0.3698±0.03 0.2987±0.02 0.3358±0.01 0.4987±0.02 0.4758±0.05 0.4983±0.01 0.3995
China 0.3676±0.02 0.3532±0.02 0.3317±0.05 0.3368±0.03 0.3670±0.03 0.4521±0.04 0.4285±0.02 0.4671±0.02 0.3880
Iran 0.3950±0.03 0.3747±0.03 0.3972±0.03 0.3384±0.04 0.3303±0.02 0.5510±0.06 0.5498±0.07 0.6086±0.03 0.4431
Greece 0.3854±0.03 0.3707±0.03 0.3461±0.02 0.3569±0.04 0.3257±0.01 0.5415±0.08 0.4459±0.02 0.5527±0.02 0.4156
Spain 0.3761±0.01 0.3386±0.02 0.3290±0.02 0.3215±0.02 0.2923±0.01 0.4630±0.03 0.4247±0.03 0.4641±0.01 0.3762
Russia 0.3686±0.01 0.3778±0.03 0.3286±0.01 0.3386±0.01 0.2912±0.02 0.4272±0.04 0.3969±0.02 0.4053±0.00 0.3668
ALL 0.3959±0.01 0.3803±0.02 0.3699±0.03 0.3276±0.02 0.3380±0.01 0.4864±0.01 0.4573±0.03 0.4981±0.01 0.4067
Corr. 0.35 0.59 0.38 0.57 0.83 0.86 0.71 0.92 -
(b) Performance results evaluated using mWS with Fasttext. Corr. represents the correlation between the FastText and BERT base embedding evaluation results.
Table 1: Automatic Evaluation: Probing performance comparison with English prompts and FmLAMA-en sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, and “b/l” stands for base/large. Bold and underline indicate the best and second-best performing cultural groups, respectively, for each model (within each column). The average Pearson correlation between mAP and mWS across LLMs is 0.72.
Bb Bl mB T5
mAP 0.60 0.49 0.48 0.56
mWS 0.48 0.38 0.36 0.41
mT5 Qwen2 Llama2 Llama3
mAP 0.38 0.05 0.04 0.13
mWS -0.06 0.05 -0.04 0.02
Table 2: Correlation between Country Dish Counts and Probing Performance.
Origin GPT-4o Human Origin GPT-4o Human Origin GPT-4o Human
Italy 52.40 55.70 U.K. 64.12 72.33 Iran 53.39 51.70
U.S. 70.06 70.26 Mexico 47.05 51.95 Greece 47.68 51.23
Turkey 55.19 53.59 India 64.22 63.72 Spain 43.02 53.16
Japan 50.76 56.07 Germany 51.71 59.02 Russia 31.42 44.86
France 52.03 58.28 China 61.52 68.96 All 56.58 -
Table 3: Manual Evaluation: Llama3 probing results evaluated using MES. The human evaluation score is the average MES from three evaluators per country group. Since evaluators vary by country group, the overall result (ALL) is shown as “-”.
Manual Evaluation Score.
MES considers only the top-$l$ predicted ingredients for evaluation in each dish, similar to the mWS metric, where $l$ is the number of reference ingredients. A predicted ingredient is considered correct if it satisfies any of the following criteria: (1) Direct Match: The ingredient is exactly in the Wikidata reference; (2) Substitutability: The ingredient can replace one in the reference; (3) Missing Traditional Ingredient: The ingredient is traditionally or commonly used in the dish, but not listed in the reference. Given a manual evaluator $\mathrm{M}\left(\cdot\right)$, the MES is defined as follows:
$\mathrm{S}_{i}$ $=$ $\frac{1}{l}\sum_{p\in\mathrm{topl}_{i}}{\mathrm{M}\left(p,\mathrm{ing}_{i}\right)}$ (9)
$\mathrm{MES}$ $=$ $\frac{1}{\left|G\right|}\sum_{i\in\mathrm{G}}{\mathrm{S}_{i}}$ (10)
where $\mathrm{M}\left(\cdot\right)$ is 1 if the top-$l$ predicted ingredient $p$ is evaluated as meeting the criteria; otherwise, 0.
5 Experiments: Existing Cultural Biases
5.1 Experimental setup
Baselines
We explore encoder-only LLMs, including BERT Devlin et al. (2019) and mBERT, encoder-decoder LLMs, such as T5 Raffel et al. (2020) and mT5 Xue et al. (2020), as well as decoder-only LLMs like Qwen2 Yang et al. (2024), Llama2 Touvron et al. (2023), and Llama3 AI@Meta (2024). Of these, BERT and T5 are English monolingual LLMs, while the remaining five are multilingual LLMs.  [^4] Monolingual LLM configurations, as well as results for the five languages on their respective filtered sub-datasets, are provided in the Appendix D.1. Additionally, we employ a ‘dumb’ baseline, in which the model is assumed to consistently predict the top-10 most common ingredients for each dish.
Cultural groups
We identify the top 30 countries with the most dishes across six languages, take their intersection to ensure sufficient data, consider geographical diversity, and ultimately narrow our focus to 14 cultural groups, as shown in Table 1.
Metrics
We evaluate all LLMs using two automated metrics. Due to cost constraints, we conduct manual evaluation only on the LLM that achieve the highest probing performance in the automated metrics. Specifically, for the evaluator $\mathrm{M}\left(\cdot\right)$ used to compute MES, we utilize GPT-4o OpenAI et al. (2024) [^5] Version: gpt-4o-2024-08-06 as a simulated evaluator and recruit real human evaluators from Prolific. [^6] https://www.prolific.co/ For each country group, we hire 3 evaluators familiar with the related cuisine, with a pay rate of £9/hour, considered mid-level on the platform. Details on the GPT-4o evaluation prompts and the human evaluation platform are provided in the Appendix C.
5.2 Automatic evaluation results
The probing results based on English prompt on the filtered dataset FmLAMA-en are illustrated in Table 1. The average Pearson correlation between mAP and mWS across all LLMs is 0.72, indicating a strong positive relationship and demonstrating that the experimental results are consistent across both automatic metrics. Overall, without fine-tuning, decoder-only LLMs exhibit significantly better cultural knowledge recall in such challenging tasks compared to encoder-only and encoder-decoder LLMs, highlighting their superior capacity as knowledge bases Petroni et al. (2019). Besides, results across all models (Avg. Column) indicate that the groups from the U.S. and India, whose official languages include English, perform the best generally. They consistently rank in the top 3 in both automatic evaluation metrics. This pattern is especially pronounced in the results for encoder-only LLMs and encoder-decoder LLMs. Apart from the Iran group [^7] See detailed analysis in Appendix E., these two cultural groups also demonstrate the best performance in decoder-only LLMs. Especially, for the ‘dumb’ baseline, the U.S. probing performance is not the best, suggesting that the data distribution does not influence the LLM’s cultural judgment capabilities. This further strengthens the persuasiveness of our experiment.
Additionally, (1) to assess the variation in model performance across different cultural groups, we calculate the Coefficient of Variation based on the absolute-match mAP metric. This serves as an indicator of the degree of cultural bias in each model’s performance. Our results show that decoder-only LLMs exhibit less cultural bias compared to other model types. Among these, the newer version, Llama3, demonstrates greater consistency in performance than Llama2. Similarly, the non-English-dominant Qwen2 LLM also shows smaller variations in cultural bias. (2) To explore the impact of different embedding methods in mWS, we use BERT Base with mean pooling to evaluate probing performance under the mWS metric. Results (see Table 8) show that the U.S. cultural group achieves the highest performance, while decoder-only LLMs exhibit significant improvement, consistent with findings using FastText. We calculate the correlation between mWS evaluations using BERT Base and FastText for each LLM. The results reveal a strong correlation for decoder-only LLMs, as shown in Table 1. (3) To explore the correlation between training data size and cultural probing performance, we use country dish counts as a proxy for Wikidata’s cultural data size and calculate their correlation with probing performance. The correlation results are show as Table 2. Based on the probing performance evaluated by the two automated metrics, LLMs like Bb, B1, and T5 consistently exhibit little correlations with cultural data size. In contrast, decoder-only LLMs show almost no correlations, indicating limited sensitivity to cultural data size as represented by dish counts. Therefore, in more advanced LLMs, cultural competency isn’t directly tied to the training sample size of specific cultural groups.
@@ 5.2 Automatic evaluation results
Figure 2: Comparative probing results on FmLAMA-en: incorporating cultural information about the origin of dishes into English prompts can enhance the probing of cultural knowledge. Significance levels are indicated by * for $p<0.1$, **∗ ∗ for $p<0.05$, and ***∗ ∗ ∗ for $p<0.01$.
5.3 Manual evaluation results
Table 1 shows that Llama3 achieve the best probing performance, so we conduct manual evaluation on its probing results. Table 3 displays the results of the GPT-4o simulator and human evaluations, showing a high Pearson correlation of 0.88. Both evaluation methods consistently show that the highest MES scores come from English-speaking cultural groups, including the U.S., U.K., and India, with the U.S. performing especially well in both manual evaluations. This aligns closely with the automatic evaluation results, reinforcing the argument that LLMs tend to exhibit cultural bias toward these groups.
Figure 3: Comparison of probing results on Llama3 with prompts in different languages, showing that the English prompt exhibits the best performance.
Figure 4: Probing results using code-switching prompts on Qwen2 and LLaMA3.
Upon reviewing the evaluation results for the GPT-4o simulator, we find that it treats synonymous expressions as Direct Matches, e.g., cooked rice vs. gohan (the Japanese term) and cauliflower vs. Brassica oleracea var. botrytis (the botanical term). Additionally, reference labels from Wikidata often exhibit inconsistencies, with terms like "shrimp or prawn" and "fish as food" marked as ingredients. Furthermore, GPT-4o flags pre-made ingredients as Substitutable; for example, flour in Garganelli with pre-made pasta. For human evaluators, evaluation results may vary due to individual subjective biases. The average Inter-Rater Reliability (IRR) across country groups, measured by Fleiss’ kappa, is 0.57, [^8] IRR and the average evaluation time for each country group are shown in Appendix C. indicating moderate agreement. While this indicates acceptable reliability for subjective tasks, it also highlights that evaluators from the same cultural background may still show variation in their judgments. Although GPT-4o can partially simulate human evaluations, its hallucination issues Ji et al. (2023) raise concerns about the reliability of its assessments. These observations underscore the inherent challenges of using automatic metrics for culturally-related tasks.
6 Analysis: Designing Better Prompts
@@
We examine the impact of different prompt settings on cross-cultural knowledge exploration, focusing on cultural background references, multilingual prompts, and code-switching scenarios.
Cultural background analysis.
@@
Using the English prompts from Figure 9 as a basis, we incorporate information about the dish country of origin into the probing prompts for each specific dish (as described in §4.2). Figure 2 presents the comparative probing results on FmLAMA-en across different LLMs, considering the inclusion of cultural information mentions in the prompts. We find that English prompts with cultural information achieve higher probing scores in capturing LLMs’ knowledge within the food domain. This suggests the importance of emphasizing the cultural background when utilizing LLMs, especially in the exploration of culture-related topics.
Multilingual prompt analysis.
We utilize prompts in six different languages to conduct knowledge probing on the multilingual LLM, Llama3.999The probing results comparison on other multilingual LLMs are shown in Appendix D.2. To ensure a fair comparison of probing results across different language prompts, we filter a sub-dataset where all food instances have both subject and ingredient labels available in all the languages involved. Figure 3 displays probing results on the filtered sub-dataset with prompts in different languages. [^10] Some Chinese and Korean objects are missing Fasttext vectors. Therefore, mWS cannot be calculated for them. The English prompt performs significantly better than prompts in other languages across all cultural groups, probably due to more training data.
Code-switching analysis.
Figure 4 presents the probing results using code-switching prompts, for each pair out of six languages, applied to the two best-performing decoder-only LLMs, Qwen2 and Llama3. When English is the main language and the dish subject is presented in another language, probing performance significantly declines but still outperforms probing conducted solely in other languages. This suggests that the instruction language is important, not just the language of the dish name. Furthermore, when the main language of the prompt is non-English and the subject language is English, performance does not significantly decline and even shows a slight improvement. This further emphasizes the dominant role of English in multilingual LLMs. The detailed probing results of each cultural group and other multilingual LLMs are in shown Appendix D.3.
Country Top 5 Most Falsely Predictions (Llama3)
Italy pasta, bread, meat, egg, pork
Japan rice, fish, soy, tofu, chicken
France butter, cream, meat, caramel, bread
Russia meat, beef, cabbage, pork, beet
(a) Coarse-grained confusions
Model Top 5 Most Frequent Predictions
BERT dried meat, ham, beef, pork, fish
T5 vinaigrette, scrambled eggs, buckwheat flour, ricotta, sultana
Llama2 rice, meat, bread, white meat, egg
Llama3 rice, meat, sugar, chicken, bread
(b) Fine-grained confusions
Table 4: Error Analysis: Coarse and fine-grained confusions in ingredient predictions by various cultural groups and models.
7 Error Analysis
A manual inspection of a data sample [^11] Conducted by one of the authors on 500 predictions by Llama3, Llama2, Bert-base-uncased and T5-base. followed by an automated evaluation on the complete data reveals recurring mistake patterns: (1) Coarse-grained confusions (foreign ingredients, e.g., coffee in a Chinese dish), where the main issue lies in the model’s lack of cultural attribute awareness for both the dish and its ingredients. (2) Fine-grained confusions (wrong local ingredients, e.g., Chinese wine in Mapo tofu), and (3) Inconsistent confusions, where the model predicts different incorrect ingredients across prompts for the same dish.
Table 4 illustrates the described error patterns. In the results for Llama3 in English, $47\%$ of the incorrect predictions for Italian dishes are the ingredient pasta, $45\%$ of the incorrect predictions for French dishes are butter, and $40\%$ of the incorrect predictions for Japanese dishes are tofu. These are examples of fine-grained confusions. Additionally, rice, wheat and flour account for $11\%$ of the Llama3’s total predictions, regardless of the dish’s origin. This is a coarse confusion, where the model often predicts the same foreign ingredient.
These errors reveal a lack of cultural knowledge about the dishes and suggest different types of guessing. While coarse and inconsistent confusions seem to reflect more general guesses, possibly influenced by the frequency of certain ingredients in the training data, fine-grained confusions indicate that the model has some knowledge of the dish’s origin and is attempting to guess a local ingredient accordingly (i.e. overfitting to the context). Specifically, for fine-grained errors, even when the cultural attributes of candidate ingredients align with the dish, LLMs still make incorrect predictions. While many studies have shown that LLMs can function as knowledge bases Nguyen et al. (2024); Pan et al. (2024) and memorize factual information Fierro et al. (2024a, b), the familiarity with specific entities is crucial in shaping their predictive expectations Du et al. (2024).